OpenStudy (anonymous):

Ryan bought a digital camera with a list price of $150 from an online store offering a 5% discount. He needs to pay $5.50 for shipping. What was Ryan's total cost

OpenStudy (anonymous):

150-5% = 142.50 + 5.50 = $148